OpenStudy (anonymous):

given that : f''(x) = x + (f'(x) /x) f'(-1)=f(1)= -1 then solve f(x)..i mean solve the differential eqn,

OpenStudy (shaik0124):

@Aryang

OpenStudy (shaik0124):

ok then it will be xf"(x)=x^2+f'(x)

OpenStudy (anonymous):

yes..then ..??

OpenStudy (anonymous):

leme try.. xf'(x) - f(x) = x^3 /3 + f(x) + C..is that right ?

OpenStudy (shaik0124):

now if u solve the equation by integration u will get answer

OpenStudy (shaik0124):

ya right@ar

OpenStudy (anonymous):

but what'd be my next step ? if i integrate now, i'll have a integral(f(x)),,wont that just get stuck ?

OpenStudy (shaik0124):

now u have to substitute the value f'(-1)=f(1)= -1

OpenStudy (anonymous):

alright,,that'll get me the value of C,,but main aim is to solve for the eqn of the curve..??

OpenStudy (anonymous):

??

OpenStudy (experimentx):

try this \[ f'(x) = u \\ f''(x) = u'\]

OpenStudy (experimentx):

there isn't f(x) ... so you can substitute f'(x) =u, this reduces to first order equation. try to solve it step wise.

OpenStudy (anonymous):

then what do i substitute for x ?

OpenStudy (experimentx):

you don't need to substitute it for x. 'u' will be function of x.

OpenStudy (anonymous):

leme try.. x u' = x^2 + u then integrate..right?

OpenStudy (experimentx):

yep ... this is linear DE. find u \[ u = f'(x) = .... \]

OpenStudy (anonymous):

okay,. u - f(x) = (x^3 /3) + f(x) then?? wont it be stuck afterwards??

OpenStudy (anonymous):

you mean the integration factor thing ?

OpenStudy (experimentx):

\[ \large u = c e^{\int {1 \over x }dx} dx + e^{\int {1 \over x }dx} \int e^{-\int {1 \over x }dx} {1 \over x} dx \]

OpenStudy (anonymous):

how'd you get that ?

OpenStudy (anonymous):

integration factor was 1/x right? you arranged it as u' - u/x = x ?

OpenStudy (experimentx):

yep

OpenStudy (anonymous):

so simplified form is u/x = x + C or u = x^2 + Cx ?

OpenStudy (anonymous):

or u = x^2 + 2x ??

OpenStudy (anonymous):

ahh..,i get it,,integrating at the next step directly solves for f(x)..thanks @experimentX

OpenStudy (experimentx):

yep \[ g\left (x, f^{n-1}x , f^{n}x \right) = 0\] solve this type of equation using this technique.

OpenStudy (anonymous):

i get it,,i know all of these,,just didnt strike me then,,thanks again..

OpenStudy (shubhamsrg):

leme try an alternative method.. d/dx ( f'(x) /x) = (x f''(x) - f'(x) )/x^2 you are given f''(x) = x + (f'(x) /x) or x f''(x) - f'(x) = x^2 thus on making this simple substitution, we have d/dx(f'(x) /x) = 1 or f'(x) = x^2 +Cx ofcorse, you can do further.. :-)

OpenStudy (anonymous):

aha,,nice..thanks!! :)